# Module schemes in invariant theory

### v1

@@ 26. Trivialization
By [37, Théorème 5.1], for any scheme $X^{\prime}$ of finite type over ${\mathbb{C}}$ such that $X^{\prime}_{\mathbb{C}}$ has a finite fundamental group, there exists a finite étale morphism $Y^{\prime}\rightarrow X^{\prime}$ such that $Y^{\prime}$ is simply connected and such that $\pi_{1}(X^{\prime})$ acts on $Y^{\prime}$ with quotient $X^{\prime}$. When $X^{\prime}\subseteq X$ is open and $X=\operatorname{Spec}R$ is a normal domain, then this can be extended to a finite morphism $Y\rightarrow X$ by going to the normalization of $X$ inside $Q(Y^{\prime})$. The group $\pi_{1}(X^{\prime})$ also acts on $Y$ with quotient $X$. It seems to be difficult to characterize, starting from $X^{\prime}=X_{\operatorname{reg}}\subseteq\operatorname{Spec}R$, whether $Y$ is an affine space. However, if we start with a group $G$ acting linearly and faithfully on ${\mathbb{A}}^{d}$ with quotient $X$, then this construction gives back ${\mathbb{A}}^{d}/G_{\beta\operatorname{refl}}$ together with the action of $G/G_{\beta\operatorname{refl}}$ on it, as was shown more generally in Lemma 22.12.
@@
(1) is clear due to the existence of the zero section. (2) The inclusion $Z_{\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}})\subseteq Z_{\operatorname{reg}}$ is an isomorphism in codimension one, due to the condition on the fibers and the normality of $X$; hence, these spaces have the same fundamental group by [44, Satz 4.B.2]. (3). The morphism $p^{-1}(X_{\operatorname{reg}})\rightarrow X_{\operatorname{reg}}$ induces a surjection $\pi_{1}(p^{-1}(X_{\operatorname{reg}}))\rightarrow\pi_{1}(X_{\operatorname{reg}})$ by part (1). The inclusion $Z_{\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}})\subseteq p^{-1}(X_{\operatorname{reg}})$ defines a surjection $\pi_{1}(Z_{\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}}))\rightarrow\pi_{1}(p^{-1}(X_{\operatorname{reg}}))$ by [44, Satz 4.B.2]. Hence, (2) gives the result. ∎
@@
Then, $\pi_{1}(U)$ is isomorphic to $\pi_{1}(Z|_{U})$ and $\pi_{1}(Z_{\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}}))$ is isomorphic to $\pi_{1}(Z_{\operatorname{reg}})$ by Lemma 26.1. The inclusions $U\subseteq X_{\operatorname{reg}}$ and $Z|_{U}\subseteq Z_{\operatorname{reg}}$ define a surjective homomorphism $\pi_{1}(Z|_{U})\rightarrow\pi_{1}(Z_{\operatorname{reg}})=H$ by [44, Satz 4.B.2]. Let $I$ be the kernel of the second homomorphism, considered as a subgroup of $\pi_{1}(U)\cong\pi_{1}(Z|_{U})$.
@@
$G=\pi_{1}(Z_{\operatorname{reg}})\cong\pi_{1}(Z_{\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}}))\longrightarrow\pi_{1}(X_{\operatorname{reg}})=G/G_{\beta\operatorname{refl}}\,.$
@@ 27. Questions
In Sections 21, 22 and 26, we worked over the complex numbers and with the topological fundamental group. What do the corresponding results in positive characteristic look like?

### v2

@@ 26. Trivialization
By [41, Théorème 5.1], for any scheme $X^{\prime}$ of finite type over ${\mathbb{C}}$ such that $X^{\prime}_{\mathbb{C}}$ has a finite fundamental group, there exists a finite étale morphism $Y^{\prime}\rightarrow X^{\prime}$ such that $Y^{\prime}$ is simply connected and such that $\pi_{1}(X^{\prime})$ acts on $Y^{\prime}$ with quotient $X^{\prime}$. When $X^{\prime}\subseteq X$ is open and $X=\operatorname{Spec}R$ is a normal domain, then this can be extended to a finite morphism $Y\rightarrow X$ by going to the normalization of $X$ inside $Q(Y^{\prime})$. The group $\pi_{1}(X^{\prime})$ also acts on $Y$ with quotient $X$. It seems to be difficult to characterize, starting from $X^{\prime}=X_{\operatorname{reg}}\subseteq\operatorname{Spec}R$, whether $Y$ is an affine space. However, if we start with a group $G$ acting linearly and faithfully on ${\mathbb{A}}^{d}$ with quotient $X$, then this construction gives back ${\mathbb{A}}^{d}/G_{\beta\operatorname{refl}}$ together with the action of $G/G_{\beta\operatorname{refl}}$ on it, as was shown more generally in Lemma 22.12.
@@
(1) is clear due to the existence of the zero section. (2) The inclusion $Z_{\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}})\subseteq Z_{\operatorname{reg}}$ is an isomorphism in codimension one, due to the condition on the fibers and the normality of $X$; hence, these spaces have the same fundamental group by [48, Satz 4.B.2]. (3). The morphism $p^{-1}(X_{\operatorname{reg}})\rightarrow X_{\operatorname{reg}}$ induces a surjection $\pi_{1}(p^{-1}(X_{\operatorname{reg}}))\rightarrow\pi_{1}(X_{\operatorname{reg}})$ by part (1). The inclusion $Z_{\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}})\subseteq p^{-1}(X_{\operatorname{reg}})$ defines a surjection $\pi_{1}(Z_{\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}}))\rightarrow\pi_{1}(p^{-1}(X_{\operatorname{reg}}))$ by [48, Satz 4.B.2]. Hence, (2) gives the result. ∎
@@
Then, $\pi_{1}(U)$ is isomorphic to $\pi_{1}(Z|_{U})$ and $\pi_{1}(Z_{\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}}))$ is isomorphic to $\pi_{1}(Z_{\operatorname{reg}})$ by Lemma 26.1. The inclusions $U\subseteq X_{\operatorname{reg}}$ and $Z|_{U}\subseteq Z_{\operatorname{reg}}$ define a surjective homomorphism $\pi_{1}(Z|_{U})\rightarrow\pi_{1}(Z_{\operatorname{reg}})=H$ by [48, Satz 4.B.2]. Let $I$ be the kernel of the second homomorphism, considered as a subgroup of $\pi_{1}(U)\cong\pi_{1}(Z|_{U})$.
@@
$G=\pi_{1}(Z_{\operatorname{reg}})\cong\pi_{1}(Z_{\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}}))\longrightarrow\pi_{1}(X_{\operatorname{reg}})=G/G_{\beta\operatorname{refl}}\,.$
@@ 27. Questions
In Sections 22 and 26, we worked over the complex numbers and with the topological fundamental group. What do the corresponding results in positive characteristic look like?